# Med-R2: Crafting Trustworthy LLM Physicians via Retrieval and Reasoning of Evidence-Based Medicine

### v4

@@ C.3 Hyper-Parameters Setting
In the fine-grained reranking phase, we treat each factor as having equal importance, hence we set the weight controlling hyper-parameter $\alpha$ in Equation 1 to 1. As for retrieval iteration settings, we performed iterative retrieval on each query within the dataset presented in Table 7 and visualized the dimensionality reduction of the retrieved document vectors across different iterations using t-SNE. Our analysis revealed that after approximately 5 iterations, the distribution of the retrieved document vectors stabilizes. Consequently, we set the maximum iterations $T$ to 5 in Algorithm 2. As depicted in Figure 9, we selected a sample query retrieval result from the MedQA-USMLE dataset to illustrate this stabilization. It is evident that after $t=5$ iterations, the vector space of the retrieved document clusters has become relatively stable. After conducting iterative retrieval for all samples in the dataset and calculating the average of the minimum distances in the retrieval space during the iterations, we obtained a value of 6.85, which we adopted as our termination alteration threshold $\delta$ in Algorithm 2.
@@
Figure 9: A case of the t-SNE visualization for retrieved documents. We visualized the projections of document embeddings onto a 2-D plane across different iterations. It is evident that aside from the significant variation in the retrieval document vector space at $t=2$, the semantics of the retrieved documents tend to stabilize in subsequent iterations. It occurs because we did not incorporate the model’s CoT sequence for the initial retrieval round ($t=1$). Instead, we began to include it starting from the second iteration onwards ($t=2$) to avoid the generation of sequences that are not only unrealistic but also detrimental to the precision of evidence document retrieval.
C.4 Additional Experimental Results: Evaluation in Real Clinical Settings
We have also evaluated Med-R2 on several benchmarks related to real-world clinical scenarios to analyze its impact in real clinical settings. For the strategy of “SFT with dataset”, we utilized the training subsets from MedQA-USMLE, MedQA-MCMLE, and MedMCQA, and followed the experimental setup outlined in Section 4.1 to fine-tune each model. Listed below are the benchmarks we employed for real-world clinical evaluation:
  • NEJMQA [^12] https://www.nejm.org/: NEJMQA is a dataset curated from The Lancet and the New England Journal of Medicine case challenges, focusing on patient diagnosis based on symptoms.
  • MedXpertQA [2]: The MedXpertQA dataset integrates specialty-specific assessments into medical benchmarking and challenging medical exam questions with real-world clinical information into medical benchmarking. Here we utilize the text part of the benchmark.
  • RareArena [^13] https://huggingface.co/datasets/THUMedInfo/RareArena: RareArena is a comprehensive rare disease diagnostic dataset with nearly 50,000 patients covering more than 4000 diseases. It includes 2 task settings, Rare Disease Screening (RDS) where the cases are truncated up to any diagnostic tests, and Rare Disease Confirmation (RDC) where the cases are truncated up to the final diagnosis.
Table 8: Comparison of Med-R2 with baselines in real clinical scenarios. The best and second best results are in bold and underlined.
Model Method NEJMQA MedXpertQA RareArena-RDC RareArena-RDS Average
Frontier Models
GPT-4o 72.20 30.37 73.76 47.50 55.96
Claude3.5-Sonnet 68.52 21.31 68.46 50.78 52.27
DeepSeek-V3 74.23 24.16 75.09 45.72 56.07
Open-Sourced Medical Models
PMC-LLaMA-7B 35.63 11.02 28.96 21.67 24.32
MEDITRON-7B 38.80 11.74 35.68 26.50 28.18
PMC-LLaMA-13B 40.60 11.44 39.86 28.98 30.22
MEDITRON-70B 66.97 18.05 74.81 46.58 51.60
Open-Sourced Base Models
Qwen2.5-7B Direct Response 41.48 11.59 32.17 23.10 27.09
Vanilla RAG 42.27 12.09 41.44 30.08 31.47
Fine-Tuning 45.11 11.67 46.39 28.85 33.01
LLM-AMT 43.80 11.94 50.15 33.18 34.77
Med-R2 45.03 12.57 55.45 38.40 38.86
LLaMA3.1-8B Direct Response 50.41 14.90 41.85 22.16 32.33
Vanilla RAG 52.01 15.50 56.23 35.06 39.70
Fine-Tuning 52.48 15.54 55.38 38.37 40.44
LLM-AMT 53.07 16.22 60.20 40.29 42.45
Med-R2 55.69 17.79 62.90 44.82 45.30
LLaMA2-13B Direct Response 45.68 11.20 38.67 24.75 30.08
Vanilla RAG 48.08 14.75 44.21 27.54 33.65
Fine-Tuning 46.51 12.17 40.09 30.66 32.36
LLM-AMT 47.66 15.69 46.45 35.93 36.43
Med-R2 50.42 14.90 52.81 40.63 39.69
Qwen2.5-14B Direct Response 45.63 11.06 43.17 26.58 31.61
Vanilla RAG 45.69 12.28 54.00 28.05 35.01
Fine-Tuning 46.67 12.30 66.32 43.66 42.24
LLM-AMT 47.44 13.73 60.03 38.81 40.00
Med-R2 48.03 14.45 72.80 46.30 45.40
Qwen2.5-32B Direct Response 43.59 15.06 59.67 31.33 37.41
Vanilla RAG 46.63 17.49 66.10 40.59 42.70
Fine-Tuning 52.58 15.71 70.20 45.08 45.89
LLM-AMT 46.82 17.90 66.75 45.40 44.22
Med-R2 50.40 18.54 72.51 55.36 49.20
LLaMA3.1-70B Direct Response 65.43 23.43 72.59 43.54 51.25
Vanilla RAG 66.40 25.95 77.70 44.88 53.73
Fine-Tuning 68.11 24.18 76.41 47.23 53.98
LLM-AMT 68.90 26.53 77.00 45.87 54.58
Med-R2 70.18 26.84 78.60 49.42 56.26
The experimental results can be found in Table 8. It is observed that in real-world clinical scenarios, where the complexity and variability of patient cases demand a higher degree of accuracy and relevance in information retrieval, the advantages of Med-R2 become more pronounced. Equipped with our Med-R2, models achieve an average improvement of 10.81% over the fine-tuning methods, but in a more cost-efficient way. Additionally, LLaMA3.1-70B + Med-R2 surpasses frontier models, including GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 in real clinical settings.
@@ Appendix D Scaling Analysis and Ablation Details
In this section, we conduct a more detailed analysis of the impact of context window size and model parameter scale on our Med-R2 . Additionally, we evaluate the contribution of each component within Med-R2 to the model’s performance in the medical domain.
D.1 Scaling for Context Window
We have selected 4 models of varying parameter scales from our main experiments to examine the impact of context window length. Considering that the training context window length for the LLaMA2 series is relatively short, at only 8K, whereas the LLaMA3.1 series models are explicitly designed to extend up to a context length of 128K. The Qwen2.5 series, trained at a 32K context length, can also be extended to 128K by modifying the max_position_embedding in the config.json file through Yarn. Consequently, we have chosen two models each from the LLaMA3.1 series and the Qwen2.5 series to conduct inference tests on context window length expansion. Experimental results are outlined in Table 9 and Figure 5.
Table 9: Scaling analysis of context window and model size for Med-R2 .
@@ D.1 Scaling for Context Window
D.2 Module Impact Analysis
We have analyzed the contributions of various modules in Med-R2 to the model’s performance in the medical domain using a default context window size of 4K, sequentially incorporating components onto the vanilla RAG framework. As shown in Table 10, we decompose the modules into three components: query reformulator, evidence reranker, and CoT generator. The optimal performance values for models with single and dual component additions are highlighted with background colors and , respectively. Overall, the evidence reranker contributed the most to models’ performances among the individual components. When combined with the query reformulator, the performance gains were even more pronounced, demonstrating a synergistic effect. The addition of the CoT generator further enhanced the model’s ability to effectively utilize retrieved medical evidence documents, providing substantial added value. However, we observed that the performance of Qwen2.5-14B on MedQA-USMLE and Qwen2.5-32B on MedQA-MCMLE was not entirely satisfactory. Upon closer examination, we found that even the base model’s direct responses were comparable to those augmented by external knowledge bases via the RAG strategy. We hypothesize that these models may have been pretrained on the MedQA dataset, thereby already possessing a robust understanding of the medical knowledge contained within such data corpus, which makes the incremental benefit from the external knowledge base less noticeable.
[…]
Qwen2.5-7B Direct Response 22.58 39.14 28.77 54.16 44.45 37.82
Vanilla RAG 58.28 64.78 32.78 55.68 54.97 53.30
[…]
+ Evidence Reranker 68.37 70.81 38.43 55.78 62.38 59.15
[…]
+ Query Reformulator, Evidence Reranker 77.85 78.67 45.62 55.99 70.74 65.77
[…]
+ Evidence Reranker, CoT Generator 75.89 73.36 42.88 55.92 68.96 63.40
Med-R2 80.67 81.23 49.18 56.74 70.96 67.76
LLaMA3.1-8B Direct Response 31.16 41.45 30.02 36.17 37.12 35.18
Vanilla RAG 55.80 59.38 35.91 47.10 43.54 48.35
+ Query Reformulator 62.47 72.85 41.66 48.72 43.98 53.94
+ Evidence Reranker 68.62 75.74 43.84 49.97 43.76 56.34
+ CoT Generator 59.87 65.63 39.45 47.71 43.61 51.25
Med-R2 77.01 84.16 52.33 52.75 45.19 62.29
LLaMA2-13B Direct Response 25.84 23.00 29.68 40.71 43.57 32.56
Vanilla RAG 36.97 22.15 35.14 52.67 45.81 38.55
+ Query Reformulator 37.32 25.64 36.98 53.83 46.67 40.09
+ Evidence Reranker 37.54 25.98 38.74 54.04 46.94 40.65
+ CoT Generator 37.44 23.12 35.93 53.22 46.03 39.15
+ Query Reformulator, Evidence Reranker 37.99 29.28 40.17 56.73 48.63 42.56
+ Query Reformulator, CoT Generator 37.58 26.88 39.65 54.94 46.98 41.21
+ Evidence Reranker, CoT Generator 38.00 27.43 41.79 56.08 47.72 42.20
Med-R2 38.06 29.34 41.96 57.12 48.89 43.07
Qwen2.5-14B Direct Response 50.01 65.23 42.85 56.93 71.60 57.32
Vanilla RAG 54.88 75.60 42.06 60.38 79.46 62.48
+ Query Reformulator 54.93 78.96 45.43 64.85 81.73 65.18
+ Evidence Reranker 55.03 78.83 45.97 65.08 81.56 65.29
[…]
Qwen2.5-32B Direct Response 16.23 87.07 66.44 68.66 80.19 63.72
Vanilla RAG 19.33 89.30 67.63 67.06 83.85 65.43
+ Query Reformulator 20.14 89.42 68.72 68.41 83.97 66.13
+ Evidence Reranker 20.83 89.88 68.88 68.47 83.99 66.41
[…]
Med-R2 24.43 90.01 69.09 68.36 84.95 67.37
LLaMA3.1-70B Direct Response 46.43 58.36 62.33 66.81 71.33 61.05
Vanilla RAG 62.66 77.91 66.63 68.78 76.95 70.59
+ Query Reformulator 70.82 80.64 69.72 71.82 73.56 73.31
+ Evidence Reranker 73.96 80.78 70.87 74.64 75.28 75.11
+ CoT Generator 68.08 78.65 68.85 70.09 73.72 71.88
Med-R2 86.37 84.58 73.36 78.24 82.82 81.07
@@ Appendix E Case Studies
Figure 10: A case study of query reformulation. We highlight the sentences within the retrieved documents that exhibit a high degree of relevance to the keywords in the query and the final answer after the query reformulation process. It is evident that prior to query reformulation (left), the retrieved evidence contains descriptions of acute liver failure mortality prognoses that are dispersed, including “hepatic encephalopathy”, “INR”, “serum bilirubin level”, etc., without emphasizing the most critical prognostic factors. Consequently, the model is unable to extract the correct answers when answering medical questions based on such documents. In contrast, after the query reformulation process (right), it can be observed that the retrieved documents consistently identify “INR” as a core prognostic feature, providing effective evidential support for the model’s responses to medical inquiries.
Figure 11: A case study of CoT generation. We highlight the keywords within the generated CoT sequences and retrieved documents that play a pivotal role in both retrieval recall and answer inference. It is observed that when the model generates CoT sequences based on the initial set of retrieved documents, it incorporates keywords such as “infections”, “autoimmune diseases”, “malignancies”, and “rheumatologic diseases”. The inclusion of these keywords enhances the comprehensiveness of the evidence documents retrieved in subsequent iterations. For instance, the second round of document retrieval encompasses causes like “systemic lupus erythematosus (SLE)” and “systemic inflammatory response syndrome (SIRS)”. Through such iterative process, the model progressively accumulates more relevant information, leading to a final answer that is notably more accurate and comprehensive compared to approaches that do not incorporate the CoT generation module.
@@ Appendix F Prompts
We present the prompts that are employed throughout our pipeline in Med-R2 . Only the English version is presented due to LaTeX compilation issues with non-English languages. Specially, prompts for query reformulation according to the Evidence-Based Medicine (EBM) categories can be found in Table 4.
@@
Prompt: Chain-of-Thought (CoT) Generator Given the provided [Context], [Question], as well as the [Retrieved Documents], please provide an answer that includes your thought process. Specifically: 1. Analyze the Question: Carefully analyze the [Question] to understand what information is being sought. 2. Review Provided Context: Examine the [Context] for any background information that can help frame the answer. 3. Consult Retrieved Documents: Go through the snippets of [Retrieved Documents] to identify sections that are directly related to [Question]. 4. Identify Key Information: Highlight the key points from the [Retrieved Documents] that address the question’s requirements. 5. Construct Thought Process: Explain how you used the information from the [Context] and the retrieved documents to form your understanding and construct your answer. 6. Provide Answer: Finally, give a clear and concise answer to the [Question], supported by the analysis of the [Retrieved Documents]. Please present your response in a way that clearly shows your reasoning and the sources of information you relied on. # Context
@@
Appendix G Ethical Considerations
Integrating Large Language Models (LLMs) into medical settings holds potential for improving healthcare but also raises ethical concerns. Med-R2 adheres to the Evidence-Based Medicine (EBM) process, balancing scientific evidence, clinical expertise, and patient values to enhance accuracy. We ensure the evidence is up-to-date and ethically sourced by continuously updating our knowledge base from reliable sources. Additionally, the retrieved knowledge corpus improves the transparency and explainability of LLMs in decision-making, making their reasoning clear and interpretable through an iterative process aligned with EBM principles.

### v5

@@ C.3. Hyper-Parameters Setting
In the fine-grained reranking phase, we treat each factor as having equal importance, hence we set the weight controlling hyper-parameter $\alpha$ in Equation 1 to 1. For retrieval iteration settings, we performed iterative retrieval on each query within the datasets in Table 7. Our analysis revealed that after approximately 5 iterations, the distribution of the retrieved document vectors stabilizes. Consequently, we set the maximum iterations $T$ to 5 in Algorithm 2. After performing iterative retrieval for all samples in the dataset and calculating the average of the minimum distances in the retrieval space during iterations, we obtained a value of 6.85, which we adopted as our termination alteration threshold $\delta$ in Algorithm 2.
@@
Figure 9. A case of the t-SNE visualization for retrieved documents. We visualized the projections of document embeddings onto a 2-D plane across different iterations. It is evident that aside from the significant variation in the retrieval document vector space at $t=2$, the semantics of the retrieved documents tend to stabilize in subsequent iterations. It occurs because we did not incorporate the model’s CoT sequence for the initial retrieval round ($t=1$). Instead, we began to include it starting from the second iteration onwards ($t=2$) to avoid the generation of sequences that are not only unrealistic but also detrimental to the precision of evidence document retrieval.
@@ Appendix D Scaling Analysis and Ablation Details
In this section, we conduct a more detailed analysis of the impact of context window size and model parameter scale on our Med-R2.
Table 8. Scaling analysis of context window and model size for Med-R2 .
@@
Scaling for Context Window  The training context windows length for the LLaMA3.1 series models are explicitly designed to extend up to a context length of 128K. The Qwen2.5 series, trained at a 32K context length, can also be extended to 128K by modifying the max_position_embedding in the config.json file through Yarn. Consequently, we have chosen two models each from the LLaMA3.1 series and the Qwen2.5 series to conduct inference tests on context window length expansion. Experimental results are outlined in Table 8 and Figure 7.
@@ Appendix E Case Studies
Figure 10. A case study of query reformulation. We highlight the sentences within the retrieved documents that exhibit a high degree of relevance to the keywords in the query and the final answer after the query reformulation process. It is evident that prior to query reformulation (left), the retrieved evidence contains descriptions of acute liver failure mortality prognoses that are dispersed, including “hepatic encephalopathy”, “INR”, “serum bilirubin level”, etc., without emphasizing the most critical prognostic factors. Consequently, the model is unable to extract the correct answers when answering medical questions based on such documents. In contrast, after the query reformulation process (right), it can be observed that the retrieved documents consistently identify “INR” as a core prognostic feature, providing effective evidential support for the model’s responses to medical inquiries.
Figure 11. A case study of CoT generation. We highlight the keywords within the generated CoT sequences and retrieved documents that play a pivotal role in both retrieval recall and answer inference. It is observed that when the model generates CoT sequences based on the initial set of retrieved documents, it incorporates keywords such as “infections”, “autoimmune diseases”, “malignancies”, and “rheumatologic diseases”. The inclusion of these keywords enhances the comprehensiveness of the evidence documents retrieved in subsequent iterations. For instance, the second round of document retrieval encompasses causes like “systemic lupus erythematosus (SLE)” and “systemic inflammatory response syndrome (SIRS)”. Through such iterative process, the model progressively accumulates more relevant information, leading to a final answer that is notably more accurate and comprehensive compared to approaches that do not incorporate the CoT generation module.
@@ Appendix F Prompts
We present the prompts employed throughout our pipeline in Med-R2 . Specially, prompts for query reformulation according to the Evidence-Based Medicine (EBM) categories can be found in Table 4.
@@
Prompt: Chain-of-Thought (CoT) Generator Given the provided [Context], [Question], as well as the [Retrieved Documents], please provide an answer that includes your thought process. Specifically: (1) Analyze the Question: Carefully analyze the [Question] to understand what information is being sought. (2) Review Provided Context: Examine the [Context] for any background information that can help frame the answer. (3) Consult Retrieved Documents: Go through the snippets of [Retrieved Documents] to identify sections that are directly related to [Question]. (4) Identify Key Information: Highlight the key points from the [Retrieved Documents] that address the question’s requirements. (5) Construct Thought Process: Explain how you used the information from the [Context] and the retrieved documents to form your understanding and construct your answer. (6) Provide Answer: Finally, give a clear and concise answer to the [Question], supported by the analysis of the [Retrieved Documents]. Please present your response in a way that clearly shows your reasoning and the sources of information you relied on. # Context